OpenStudy (anonymous):

Which of the following is not equivalent to the formula for the circumference of a circle, C = 2 pi r ? 2 times pi over c equals r c over quantity 2 times r equals pi c over quantity r times pi equals 2 c over quantity 2 times pi equals r

OpenStudy (mysesshou):

2 times pi over c equals r \[2*pi/C=r\] x c over quantity 2 times r equals pi \[C/2r = pi\] c over quantity r times pi equals 2 \[C/r*pi = 2\] c over quantity 2 times pi equals r \[C/2*pi=r\] first one is not equivalent.

OpenStudy (anonymous):

thanks